# Deepfake Generation and Detection: A Benchmark and Survey

### v3

@@ 3.1.1 Face Swapping
In this section, we review face swapping methods from the perspective of basic architecture, which can be mainly divided into four categories and summarized in Tab. II.
$\bullet$ Traditional Graphics. As representative early implementations, traditional graphics methods in the implementation path can be divided into two categories: 1) Key information matching and fusion. Methods [174, 196, 178] grounded in critical information matching and fusion are geared towards substituting corresponding regions by aligning key points within facial regions of interest (ROIs), such as the mouth, eyes, nose, and mouth, between the source and target images. Following this, additional procedures such as boundary blending and lighting adjustments are executed to produce the resulting image. Bitouk et al[174] accomplish automated face replacement by constructing a substantial face database to locate faces with akin poses and lighting conditions for substitution. Meanwhile, Nirkin et al[178] enhance keypoint matching and segmentation accuracy by incorporating a Fully Convolutional Network (FCN) into their method. 2) The construction of a 3D prior model for facial parameterization. Methods [173, 175] based on constructing a 3D prior and introducing a facial parameter model often involve building a facial parameter model using 3DMM technology based on a pre-collected face database. After matching the facial information of the source image with the constructed face model, specific modifications are made to the relevant parameters of the facial parameter model to generate a completely new face. Dale et al[175] utilize 3DMM to track facial expressions in two videos, enabling face swapping in videos. Some methods [176, 197] have explored scenarios involving significant pose differences between the source and target images. Lin et al[176] constructs a 3D face model from frontal faces, renderable in any pose. Guo et al[197] utilize plane parameterization and affine transformation to establish a one-to-one dense mapping between 2D graphics. Traditional computer graphics methods solve basic face-swapping problems, exploring full automation to enhance generalization. However, these methods are constrained by the need for similarities in pose and lighting between source and target images. They also face challenges like low image resolution, modification of target attributes, and poor performance in extreme lighting and occlusion scenarios.
@@
1) Early GAN-based methods [48, 198, 199, 183] address issues related to the similarity of pose and lighting between source and target images. DepthNets [198] combines GANs with 3DMM to map the source face to any target geometry, not limited to the geometric shape of the target template. This allows it to be less affected by differences in pose between the source and target faces. However, they face challenges in generalizing the trained model to unknown faces.
@@
2) Improved Generalizability. To improve the model’s generalization, many efforts [118, 180, 184] are made to explore solutions. Combining GANs with VAEs, the model [200, 117] encodes and processes different facial regions separately. FSGAN [180] integrates face reenactment with face swap, designing a facial blending network to mix two faces seamlessly. SimSwap [184] introduces an identity injection module to avoid integrating identity information into the decoder. However, these methods suffer from low resolution and significant attribute loss and need help to handle facial occlusions effectively.
@@
3) Resolution Upgrading. Some methods [185, 201, 202] provide solutions to enhance the resolution of generated images. MegaFS [185] introduces the first single-lens face swapping method at the million-pixel level. The face encoder no longer compresses facial information but represents it in layers, achieving more detailed preservation. StyleIPSB [202] constrains semantic attribute codes within the subspace of StyleGAN, thereby fixing certain semantic information during face swapping to preserve pore-level details.
@@
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods [203, 204, 189, 205] introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace [203] introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace [189] introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ [205] improves upon FlowFace by utilizing a pre-trained Mask Autoencoder to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
@@
5) Facial Masking Artifacts. Some methods [182, 180, 206, 207] have partially alleviated the artifacts caused by facial occlusion. FSGAN [180] designs a restoration network to estimate missing pixels. E4S [206] redefines the face-swapping problem as a mask-exchanging problem for specific information. It utilizes a mask-guided injection module to perform face swapping in the latent space of StyleGAN. However, overall, the methods above have not thoroughly addressed the issue of artifacts in generated images under extreme occlusion conditions.
@@
6) Trade-offs between Identity-replacement and Attribute-retention. In addition to the occlusion issues that need further handling, researchers [208, 13] discover that the balance between identity replacement and attribute preservation in generated images seems akin to a seesaw. Many methods [209, 188, 210] explore the equilibrium between identity replacement and attribute retention. InfoSwap [208] aims to decouple identity and attribute information in face swapping by leveraging the information bottleneck theory. It seeks controlled swapping of identity between source and target faces. StyleSwap [209] introduces a novel swapping guidance strategy, the ID reversal, to enhance the similarity of facial identity in the output. Shiohara et al[13] propose BlendFace, using an identity encoder that extracts identity features from the source image and uses it as identity distance loss, guiding the generator to produce facial exchange results.
@@
7) Model Light-weighting is also an important topic with profound implications for the widespread application of models. FastSwap [211] achieves this by innovating a decoder block called Triple Adaptive Normalization (TAN), effectively integrating identity information from the source image and pose information from the target image. XimSwap [12] modifies the design of convolutional blocks and the identity injection mechanism, successfully deploying on STM32H743.
$\bullet$ Diffusion-based. The latest studies [191, 192, 114, 9] in this area produce promising generation results. DiffSwaps [192] redefines the face swapping problem as a conditional inpainting task. Liu et al[9] introduce a multi-modal face generation framework and achieved this by introducing components such as balanced identity and expression encoders to the conditional diffusion model, striking a balance between identity replacement and attribute preservation during the generation process. As a novel facial generalist model, FaceX [114] can achieve various facial tasks, including face swapping and editing. Leveraging the pre-trained StableDiffusion [7] has significantly improved the quality and model training speed.
$\bullet$ Alternative Techniques. There are also some methods that stand independently from the above classifications that are discussed here collectively. Fast Face-swap [212] views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods [193, 194] apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer [57], TransFS [194] obtains rich facial features, enabling facial swapping in high-resolution images. ReliableSwap [213] enhances the model’s identity preservation capabilities by constructing a reliable supervisor called the "cyclic triplet." However, it has limitations in preserving attribute information.
@@
TABLE III: Overview of representative face reenactment methods. Notations: ➊ Voxceleb, ➋ Self-build, ➌ Voxceleb2, ➍ TalkingHead-1KH, ➎ CelebV-HQ, ➏ VFHQ, ➐ RaFD, ➑ VGGFace, ➒ CelebV, ➓FFHQ. Abbreviations: Expression (Exp).
@@
Face2Face [214] CVPR’16 Exp, Pose Using 3D facial reconstruction, transfer facial expressions and pose from a driving character to a target character via affine transforma -tion, then generate the final video through rendering techniques.
Kim et al[215] TOG’18 Exp, Pose, Blink Using synthesized rendering images of a parameterized face model as input, creating lifelike video frames for the target actor.
Kim et al[216] TOG’19 Lip,Exp, Pose Built on a recurrent generative adversarial network, it employs a hierarchical neural face renderer to synthesize realistic video frames.
Head2Head [14] FG’21 Exp, Pose, Blink, Gaze The model consists of two stages: facial reconstruction and tracking in the first stage, followed by video rendering in the second stage.
HeadGAN [217] ECCV’21 Exp, Pose Using 3DMM for facial modeling provides a 3D prior to the GAN, effectively guiding the generator to accurately recover pose and expression from the target frame.
Face2Faceρ [218] ECCV’22 Exp, Pose Decoupling the actor’s facial appearance and motion information with two separate encodings allows the network to learn facial app -earance and motion priors.
PECHead [219] CVPR’23 Lip, Exp, Pose ➌➍➎➏ A novel multi-scale feature alignment module for motion perception is proposed to minimize distortion during motion transmission.
@@
X2Face [220] CVPR’18 Exp, Pose The training is realized in two stages: the first stage guides the generation frames towards the driving frames and the second stage accomplishes the preservation of the source identity information through the identity loss function.
Zakharov et al[221] ICCV’19 Exp, Pose ➊➌ Proposed a meta-learning framework for adversarial generative models, reducing the required training data size.
FReeNet [222] CVPR’20 Exp, Pose [85] A new triple perceptual loss is proposed to richly reproduce facial details of the face.
Zakharov et al[223] ECCV’20 Exp, Pose Decomposed facial information into two layers for modeling: the first layer synthesizes coarse images related to the pose using a small neural network, and the second layer defines texture images unrelated to the pose, containing high-frequency details.
MarioNETte [224] AAAI’20 Exp, Pose ➊➒ Introduced image attention blocks, target feature alignment modules, and landmark transformers, enhancing the model’s performance when generalizing to unknown individual identities.
DG [15] CVPR’22 Exp, Pose ➊➌➐ A proposed dual generator model network for large pose face reproduction.
Doukas et al[225] TPAMI’23 Exp, Pose, Gaze [226][227] Eye gaze control in the generated video is implemented to further enhance visual realism.
MetaPortrait [228] CVPR’23 Exp, Pose By establishing dense facial keypoint matching, accurate deformation field prediction is achieved, and the model training is expedited based on the meta-learning philosophy.
Yang et al[229] AAAI’24 Exp, Pose ➊➌➍ The facial tri-plane is represented by canonical tri-plane, identity deformation, and motion components, achieving face reenactment without the need for 3D parameter model priors.
@@
HiDe-NeRF [230] CVPR’23 Lip, Exp, Pose ➊➌➍ High-fidelity and free-viewing talking head synthesis using deformable neural radiation fields.
HyperReenact [16] ICCV’23 Exp, Pose ➊➌ Exploiting the effectiveness of hypernetworks in real image inversion tasks and extending them to real image manipulation.
Stylemask [231] FG’23 Exp, Pose This work optimizes a Mask Network and combines it with StyleGAN2’s style potential space S in order to achieve the separation of facial pose and expression of the target image from the identity features of the source image.
Bounareli et al[232] IJCV’24 Exp, Pose ➊➌ In GANs’ latent space, head pose and expression changes are decoupled, achieving near-real outputs through real image embedding.
@@
ICface [233] WACV’20 Exp, Pose The model is decoupled and driven by interpretable control signals that can be obtained from multiple sources such as external driving videos and manual controls.
Oorloff et al[234] ICCV’23 Lip, Exp, Pose Identity and attribute decomposition are realized in StyleGAN2’s latent space, and a cyclic manifold adjustment technique enhances facial reconstruction results.
Xue et al. [235] TOMM’23 Exp, Pose ➊➌ High-fidelity facial generation is achieved by using information-rich Projected Normalized Coordinate Code (PNCC) and eye maps, replacing sparse facial landmark representations.
@@ 3.1.2 Face Reenactment
This section reviews current methods from four points: 3DMM-based, landmark matching, face feature decoupling, and self-supervised learning. We summarize them in Tab. III.
@@
$\bullet$ 3DMM-based. Some methods [14, 218, 219] utilize 3DMM to construct a facial parameter model as an intermediary for transferring information between the source and target. In particular, Face2Faceρ [218], based on 3DMM, consists of a u-shaped rendering network driven by head pose and facial motion fields and a hierarchical coarse-to-fine motion network guided by landmarks at different scales. However, some methods [214, 215, 216] exhibit visible artifacts in the background when dealing with significant head movements in the input images. To address issues such as incomplete attribute decoupling in facial reproduction tasks, PECHead [219] models facial expressions and pose movements. It combines self-supervised learning of landmarks with 3D facial landmarks and introduces a new motion-aware multi-scale feature alignment module to eliminate artifacts that may arise from facial motion.
@@
$\bullet$ Landmark Matching. This kind of methods [220, 223, 224, 15, 225] aim to establish a mapping relationship between semantic objects in the facial regions of the driving source and the target source through landmarks. Based on this mapping relationship [221, 222], the transfer of facial movement information is achieved. In particular, X2Face [220] achieves one-to-one driving of facial expressions and poses for each frame of the character in the driving video and source video. To address the challenge of reproducing large head poses in facial reenactment, Xu et al[15] propose a dual-generator network incorporating a 3D landmark detector into the model. Free-headgan [225] comprises a 3D keypoint estimator, an eye gaze estimator, and a generator built on the HeadGAN architecture. The 3D keypoint estimator addresses the regression of deformations related to 3D poses and expressions. The eye gaze estimator controls eye movement in videos, providing finer details. MetaPortrait [228] achieves accurate distortion field prediction through dense facial keypoint matching and accelerates model training based on meta-learning principles, delivering excellent results on limited datasets.
@@
$\bullet$ Feature Decoupling. The latent feature decoupling and driving methods [16, 231, 230, 232, 229] aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact [16] uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask [231] separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF [230] employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
@@
$\bullet$ Self-supervised Learning. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels [233, 236, 234, 237]. Oorloff et al[234] employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et al[237] utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide motion field estimation, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
@@
TABLE IV: Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ. Full table can be viewed in Appendix B.
@@
Audio / Text - Driven Chen et al[120] CVPR’19 Inability to control pose and emotional variations. Proposed a cascaded approach, using facial landmarks as an intermediate high-level representation.
@@
Wav2Lip [106] ICMR’20 ➊➎➐ Poor resolution, inability to control pose and emotional. A new evaluation framework and a dataset for training mouth synchronization are proposed.
@@
MakeItTalk [238] TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
@@
Ji et al[239] CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
@@
Gan et al[240] ICCV’23 ➊➋➌ Insufficient emotional output diversity. Proposed a two-stage architecture, implementing speaker generation independent of emotion and emb -edding emotion information.
@@
SadTalker [241] CVPR’23 ➏➒ Lack emotional and other latent attributes control. Based on the idea of 3DMM and conditional VAE, 3D coefficients controlling facial motion and expre -ssion are generated from audio to realize the reproduction of accurate faces from audio.
@@
EmoTalk [242] ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
@@
TalkLip [243] CVPR’23 ➊➎ Inability to control pose and emotional variations. Pre-trained lip-reading experts are employed to penalize incorrect lip-reading predictions in the synth -esized videos.

### v4

@@ 3.1.1 Face Swapping
In this section, we review face swapping methods from the perspective of basic architecture, which can be mainly divided into four categories and summarized in  Table 2.
$\bullet$ Traditional Graphics. As representative early implementations, traditional graphics methods in the implementation path can be divided into two categories: 1) Key information matching and fusion. Methods bitouk2008face ; sunkavalli2010multi ; nirkin2018face ground in critical information matching and fusion are geared towards substituting corresponding regions by aligning key points within facial regions of interest (ROIs), such as the mouth, eyes, nose, and mouth, between the source and target images. Following this, additional procedures such as boundary blending and lighting adjustments are executed to produce the resulting image. Bitouk et albitouk2008face accomplish automated face replacement by constructing a substantial face database to locate faces with akin poses and lighting conditions for substitution. Meanwhile, Nirkin et alnirkin2018face enhance keypoint matching and segmentation accuracy by incorporating a Fully Convolutional Network (FCN) into their method. 2) The construction of a 3D prior model for facial parameterization. Methods blanz2004exchanging ; dale2011video based on constructing a 3D prior and introducing a facial parameter model often involve building a facial parameter model using 3DMM technology based on a pre-collected face database. After matching the facial information of the source image with the constructed face model, specific modifications are made to the relevant parameters of the facial parameter model to generate a completely new face. Dale et aldale2011video utilize 3DMM to track facial expressions in two videos, enabling face swapping in videos. Some methods lin2012face ; guo2019face explore scenarios involving significant pose differences between the source and target images. Lin et allin2012face construct a 3D face model from frontal faces, renderable in any pose. Guo et alguo2019face utilize plane parameterization and affine transformation to establish a one-to-one dense mapping between 2D graphics. Traditional computer graphics methods solve basic face-swapping problems, exploring full automation to enhance generalization. However, these methods are constrained by the need for similarities in pose and lighting between source and target images. They also face challenges like low image resolution, modification of target attributes, and poor performance in extreme lighting and occlusion scenarios.
@@
1) Early GAN-based methods bao2017cvae ; moniz2018unsupervised ; liu2023deepfacelab ; zhu2020deepfakes address issues related to the similarity of pose and lighting between source and target images. DepthNets moniz2018unsupervised combines GANs with 3DMM to map the source face to any target geometry, not limited to the geometric shape of the target template. This allows it to be less affected by differences in pose between the source and target faces. However, they face challenges in generalizing the trained model to unknown faces.
@@
2) Improved Generalizability. To improve the model’s generalization, many efforts sun2018hybrid ; nirkin2019fsgan ; chen2020simswap are made to explore solutions. Combining GANs with VAEs, the model bao2018towards ; natsume2018rsgan encodes and processes different facial regions separately. FSGAN nirkin2019fsgan integrates face reenactment with face swap, designing a facial blending network to mix two faces seamlessly. SimSwap chen2020simswap introduces an identity injection module to avoid integrating identity information into the decoder. However, these methods suffer from low resolution and significant attribute loss and need help to handle facial occlusions effectively.
@@
3) Resolution Upgrading. Some methods zhu2021one ; xu2022designing ; jiang2023styleipsb provide solutions to enhance the resolution of generated images. MegaFS zhu2021one introduces the first single-lens face swapping method at the million-pixel level. The face encoder no longer compresses facial information but represents it in layers, achieving more detailed preservation. StyleIPSB jiang2023styleipsb constrains semantic attribute codes within the subspace of StyleGAN, thereby fixing certain semantic information during face swapping to preserve pore-level details.
@@
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods wang2021hififace ; ren2023reinforced ; zeng2023flowface ; zhang2023flowface++ introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace wang2021hififace introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace zeng2023flowface introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ zhang2023flowface++ improves upon FlowFace by utilizing a pre-trained Mask Autoencoder to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
@@
5) Facial Masking Artifacts. Some methods li2019faceshifter ; nirkin2019fsgan ; liu2023fine ; rosberg2023facedancer have partially alleviated the artifacts caused by facial occlusion. FSGAN nirkin2019fsgan designs a restoration network to estimate missing pixels. E4S liu2023fine redefines the face-swapping problem as a mask-exchanging problem for specific information. It utilizes a mask-guided injection module to perform face swapping in the latent space of StyleGAN. However, overall, the methods above have not thoroughly addressed the issue of artifacts in generated images under extreme occlusion conditions.
@@
6) Trade-offs between Identity-replacement and Attribute-retention. In addition to the occlusion issues that need further handling, researchers gao2021information ; shiohara2023blendface discover that the balance between identity replacement and attribute preservation in generated images seems akin to a seesaw. Many methods xu2022styleswap ; kim2022smooth ; liu2023high explore the equilibrium between identity replacement and attribute retention. InfoSwap gao2021information aims to decouple identity and attribute information in face swapping by leveraging the information bottleneck theory. It seeks controlled swapping of identity between source and target faces. StyleSwap xu2022styleswap introduces a novel swapping guidance strategy, the ID reversal, to enhance the similarity of facial identity in the output. Shiohara et alshiohara2023blendface propose BlendFace, using an identity encoder that extracts identity features from the source image and uses it as identity distance loss, guiding the generator to produce facial exchange results.
@@
7) Model Light-weighting is also an important topic with profound implications for the widespread application of models. FastSwap yoo2023fastswap achieves this by innovating a decoder block called Triple Adaptive Normalization (TAN), effectively integrating identity information from the source image and pose information from the target image. XimSwap ancilotto2023ximswap modifies the design of convolutional blocks and the identity injection mechanism, successfully deploying on STM32H743.
@@
Table 3: Overview of representative face reenactment methods. Notations: ➊ Voxceleb, ➋ Self-build, ➌ Voxceleb2, ➍ TalkingHead-1KH, ➎ CelebV-HQ, ➏ VFHQ, ➐ RaFD, ➑ VGGFace, ➒ CelebV, ➓FFHQ. Abbreviations: Expression (Exp).
@@
Face2Face thies2016face2face CVPR’16 Exp, Pose Using 3D facial reconstruction, transfer facial expressions and pose from a driving character to a target character via affine transforma -tion, then generate the final video through rendering techniques.
Kim et alkim2018deep TOG’18 Exp, Pose, Blink Using synthesized rendering images of a parameterized face model as input, creating lifelike video frames for the target actor.
Kim et alkim2019neural TOG’19 Lip,Exp, Pose Built on a recurrent generative adversarial network, it employs a hierarchical neural face renderer to synthesize realistic video frames.
Head2Head koujan2020head2head FG’21 Exp, Pose, Blink, Gaze The model consists of two stages: facial reconstruction and tracking in the first stage, followed by video rendering in the second stage.
HeadGAN doukas2021headgan ECCV’21 Exp, Pose Using 3DMM for facial modeling provides a 3D prior to the GAN, effectively guiding the generator to accurately recover pose and expression from the target frame.
Face2Faceρ yang2022face2face ECCV’22 Exp, Pose Decoupling the actor’s facial appearance and motion information with two separate encodings allows the network to learn facial app -earance and motion priors.
PECHead gao2023high CVPR’23 Lip, Exp, Pose ➌➍➎➏ A novel multi-scale feature alignment module for motion perception is proposed to minimize distortion during motion transmission.
@@
X2Face wiles2018x2face CVPR’18 Exp, Pose The training is realized in two stages: the first stage guides the generation frames towards the driving frames and the second stage accomplishes the preservation of the source identity information through the identity loss function.
Zakharov et alzakharov2019few ICCV’19 Exp, Pose ➊➌ Proposed a meta-learning framework for adversarial generative models, reducing the required training data size.
FReeNet zhang2020freenet CVPR’20 Exp, Pose moore2010multi A new triple perceptual loss is proposed to richly reproduce facial details of the face.
Zakharov et alzakharov2020fast ECCV’20 Exp, Pose Decomposed facial information into two layers for modeling: the first layer synthesizes coarse images related to the pose using a small neural network, and the second layer defines texture images unrelated to the pose, containing high-frequency details.
MarioNETte ha2020marionette AAAI’20 Exp, Pose ➊➒ Introduced image attention blocks, target feature alignment modules, and landmark transformers, enhancing the model’s performance when generalizing to unknown individual identities.
DG hsu2022dual CVPR’22 Exp, Pose ➊➌➐ A proposed dual generator model network for large pose face reproduction.
Doukas et aldoukas2023free TPAMI’23 Exp, Pose, Gaze kellnhofer2019gaze360 zhang2015appearance Eye gaze control in the generated video is implemented to further enhance visual realism.
MetaPortrait zhang2023metaportrait CVPR’23 Exp, Pose By establishing dense facial keypoint matching, accurate deformation field prediction is achieved, and the model training is expedited based on the meta-learning philosophy.
Yang et alyang2023learning AAAI’24 Exp, Pose ➊➌➍ The facial tri-plane is represented by canonical tri-plane, identity deformation, and motion components, achieving face reenactment without the need for 3D parameter model priors.
FSRTrochow2024fsrt CVPR’24 Exp, Pose The Transformer-based encoder-decoder effectively encodes attributes and improves action transmission quality.
@@
HiDe-NeRF li2023one CVPR’23 Lip, Exp, Pose ➊➌➍ High-fidelity and free-viewing talking head synthesis using deformable neural radiation fields.
HyperReenact bounareli2023hyperreenact ICCV’23 Exp, Pose ➊➌ Exploiting the effectiveness of hypernetworks in real image inversion tasks and extending them to real image manipulation.
Stylemask bounareli2023stylemask FG’23 Exp, Pose This work optimizes a Mask Network and combines it with StyleGAN2’s style potential space S in order to achieve the separation of facial pose and expression of the target image from the identity features of the source image.
Bounareli et albounareli2024one IJCV’24 Exp, Pose ➊➌ In GANs’ latent space, head pose and expression changes are decoupled, achieving near-real outputs through real image embedding.
@@
ICface tripathy2020icface WACV’20 Exp, Pose The model is decoupled and driven by interpretable control signals that can be obtained from multiple sources such as external driving videos and manual controls.
Oorloff et aloorloff2023robust ICCV’23 Lip, Exp, Pose Identity and attribute decomposition are realized in StyleGAN2’s latent space, and a cyclic manifold adjustment technique enhances facial reconstruction results.
Xue et al. xue2023high TOMM’23 Exp, Pose ➊➌ High-fidelity facial generation is achieved by using information-rich Projected Normalized Coordinate Code (PNCC) and eye maps, replacing sparse facial landmark representations.
$\bullet$ Diffusion-based. The latest studies kim2022diffface ; zhao2023diffswap ; han2023generalist ; liu2024towards in this area produce promising generation results. DiffSwaps zhao2023diffswap redefines the face swapping problem as a conditional inpainting task. Liu et alliu2024towards introduce a multi-modal face generation framework and achieved this by introducing components such as balanced identity and expression encoders to the conditional diffusion model, striking a balance between identity replacement and attribute preservation during the generation process. As a novel facial generalist model, FaceX han2023generalist can achieve various facial tasks, including face swapping and editing. Leveraging the pre-trained StableDiffusion blattmann2023stable has significantly improved the quality and model training speed.
$\bullet$ Alternative Techniques. Some methods stand independently from the above classifications that are discussed here collectively. Fast Face-swap korshunova2017fast views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods cui2023face ; cao2023transfs apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer liu2021swin , TransFS cao2023transfs obtains rich facial features, enabling facial swapping in high-resolution images. ReliableSwap yuan2023reliableswap enhances the model’s identity preservation capabilities by constructing a reliable supervisor called the "cyclic triplet." However, it has limitations in preserving attribute information.
@@ 3.1.2 Face Reenactment
This section reviews current methods from four points: 3DMM-based, landmark matching, face feature decoupling, and self-supervised learning. We summarize them in Table 3.
@@
$\bullet$ 3DMM-based. Some methods koujan2020head2head ; yang2022face2face ; gao2023high utilize 3DMM to construct a facial parameter model as an intermediary for transferring information between the source and target. In particular, Face2Faceρ yang2022face2face , based on 3DMM, consists of a u-shaped rendering network driven by head pose and facial motion fields and a hierarchical coarse-to-fine motion network guided by landmarks at different scales. However, some methods thies2016face2face ; kim2018deep ; kim2019neural exhibit visible artifacts in the background when dealing with significant head movements in the input images. To address issues such as incomplete attribute decoupling in facial reproduction tasks, PECHead gao2023high models facial expressions and pose movements. It combines self-supervised learning of landmarks with 3D facial landmarks and introduces a new motion-aware multi-scale feature alignment module to eliminate artifacts that may arise from facial motion.
@@
$\bullet$ Landmark Matching. This kind of methods wiles2018x2face ; zakharov2020fast ; ha2020marionette ; hsu2022dual ; doukas2023free aim to establish a mapping relationship between semantic objects in the facial regions of the driving source and the target source through landmarks. Based on this mapping relationship zakharov2019few ; zhang2020freenet , the transfer of facial movement information is achieved. In particular, X2Face wiles2018x2face achieves one-to-one driving of facial expressions and poses for each frame of the character in the driving video and source video. To address the challenge of reproducing large head poses in facial reenactment, Xu et alhsu2022dual propose a dual-generator network incorporating a 3D landmark detector into the model. Free-headgan doukas2023free comprises a 3D keypoint estimator, an eye gaze estimator, and a generator built on the HeadGAN architecture. The 3D keypoint estimator addresses the regression of deformations related to 3D poses and expressions. The eye gaze estimator controls eye movement in videos, providing finer details. MetaPortrait zhang2023metaportrait achieves accurate distortion field prediction through dense facial keypoint matching and accelerates model training based on meta-learning principles, delivering excellent results on limited datasets.
@@
$\bullet$ Feature Decoupling. The latent feature decoupling and driving methods bounareli2023hyperreenact ; bounareli2023stylemask ; li2023one ; bounareli2024one ; yang2023learning aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact bounareli2023hyperreenact uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask bounareli2023stylemask separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF li2023one employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
@@
$\bullet$ Self-supervised Learning. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels tripathy2020icface ; zeng2020realistic ; oorloff2023robust ; zhang2023exploiting . Oorloff et aloorloff2023robust employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et alzhang2023exploiting utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide motion field estimation, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
@@
Table 4: Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ.
@@
Audio / Text - Driven Chen et alchen2018lip ECCV’18 cooke2006audio 0Audiovisual Poor resolution, inability to control pose and emotional. Proposed a novel generator network and a comprehensive model with four complementary losses, as well as a new audio-visual related loss function to guide video generation.
Zhou et alzhou2019talking AAAI’19 Inability to control pose and emotional variations. Generate high-quality talking face videos by disentangling audio-visual representations.
@@
Chen et alchen2019hierarchical CVPR’19 Inability to control pose and emotional variations. Proposed a cascaded approach, using facial landmarks as an intermediate high-level representation.
@@
Wav2Lip prajwal2020lip ICMR’20 ➊➎➐ Poor resolution, inability to control pose and emotional. A new evaluation framework and a dataset for training mouth synchronization are proposed.
@@
MakeItTalk zhou2020makelttalk TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
@@
Ji et alji2021audio CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
SPACE gururani2023space ICCV’23 ➋➌ Lack emotional and other latent attributes control Constructed a novel facial intermediate representation, achieving control overhead pose, blinking, and gaze direction.
@@
Gan et algan2023efficient ICCV’23 ➊➋➌ Insufficient emotional output diversity. Proposed a two-stage architecture, implementing speaker generation independent of emotion and emb -edding emotion information.
@@
SadTalker zhang2023sadtalker CVPR’23 ➏➒ Lack emotional and other latent attributes control. Based on the idea of 3DMM and conditional VAE, 3D coefficients controlling facial motion and expre -ssion are generated from audio to realize the reproduction of accurate faces from audio.
Zhongetczhong2023identity CVPR’23 ➎➐ afouras2018deep Lack emotional and other latent attributes control. A two-stage framework for step-by-step audio-to-face action generation is proposed.
@@
EmoTalk peng2023emotalk ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
@@
TalkLip wang2023seeing CVPR’23 ➊➎ Inability to control pose and emotional variations. Pre-trained lip-reading experts are employed to penalize incorrect lip-reading predictions in the synth -esized videos.